OpenStudy (anonymous):

simplify the expression: cotx+1/cscx

OpenStudy (anonymous):

is it [cot(x) +1] / csc(x) or cot(x) + (1/csc(x))?

OpenStudy (anonymous):

the first way u have it. and i know to separate them, but then i get confused after that

OpenStudy (anonymous):

\[\cot(x) = \frac{\cos(x)}{\sin(x)}\] \[\csc(x) = \frac{1}{\sin(x)}\] so \[\frac{\frac{\cos(x)}{\sin(x)}+1}{\csc(x)}\] \[\frac{\frac{\cos(x)}{\sin(x)}+1}{\frac{1}{\sin(x)}}\] if you multiply the top and bottom by sin(x) then you get \[\cos(x) + \sin(x)\]